OpenStudy (wolframwizard):

Find the sum of the infinite series: (1/2)-(1/2)-(1/3)+(1/3)-(1/16)-(3/8)-(1/6)+(1/12)+(1/128)-(9/32)-(1/12)+(1/48)...

OpenStudy (evoker):

What is the pattern here?, i'm not seeing anything immediately.

OpenStudy (mww):

yeah initially terms cancel but after the pattern is not very clearcut.

OpenStudy (wolframwizard):

I'm not sure what the pattern is

OpenStudy (wolframwizard):

The answer is -16/9 but I need to know how to solve the problem

jhonyy9 (jhonyy9):

- the 4 terms cancel -(1/16)-(3/8)= -(7/16) -(1/6)+(1/12)= -(1/12) +(1/128)-(9/32)= -(35/128) -(1/12)+(1/48)= -(3/48) = -(1/16) assume to -(7/16) from the top this -(1/16) -(7/16) = -(8/16) = -(1/2) assume with second line result -(1/12) -(1/12)-(1/2) = -(7/12) do you can writing the next terms again ?

jhonyy9 (jhonyy9):

Loser any idea about this ?

OpenStudy (loser66):

No, I don't have any idea :(